7. A 65 kg trampoline artist jumps vertically upward from the top of platform with speed of
5.0 m/s .
a) How fast is he going as he lands on the trampoline 3.0 m below?
b) If the trampoline behaves like a spring with spring stiffness constant 6.2 x 10 4 N/m.
How far does he depress it?
1
Expert's answer
2020-05-21T13:53:20-0400

a)

applying "v^2=u^2+2gS"

"v^2=25+2\times10\times3=95\\v=9.75m/s"


b)

if the trampoline behaves like a spring then the depression in the spring in the spring will create potential energy

so applying energy conservation

"\frac12mv^2=\frac12kx^2"

"65\times(9.75)^2=6.2\times10^4\times x^2"

"x=31.56\times10^{-2}m"

x=31.56cm